$@MID@$: 277 Date: Sat, 07 Dec 2019 19:16:01 +0000 Author: Daniel Mazin, Koji Noda Type: Hardware Category: Hardware Subject: ALS East: problem and preliminary solution Attachment: 191207_192642_a6cb1e40-f3ab-4f8c-8705-3f34ddaba813.JPG,191207_192825_IMG_2040.jpg,191207_192904_IMG_2042.jpg,191207_192913_IMG_2044.jpg,191207_192923_IMG_2045.jpg,191207_193028_Screenshot_2019-12-07_at_19.22.57.png Encoding: HTML ========================================
On Dec 6, Dario, Frederic and Yuji showed to Koji Noda and Daniel Mazin how to operate the new ALS. The idea was to do an unlock and lock again, as the weather was still bad. The automatic part went more or less OK ending in error state (excpected). We then went to EAST ALS to open the main actuator manually. After moving the actuator manually, the status of the ALS was either locked not unlocked. We realized that secondary sensor for South part was reacting incorrectly (close to metal OFF, far from metal ON), and we thought it is the reason. We removed the endswictch screw but the status did not change. After pressing again Unlock we realized that the secondary South plier squeezed the sensor cable and destroyed it (see Attachment 1). We consulted Ino and closed the ALS EAST manually (all three actuators) because the eather was getting worse.
Next day, Dec 7, we (Daniel and Koji) examined the sensor cable and confirmed it is completely distroyed. We introduced a patch cable and few adapters as a temporary repair (See attachements 2,3,4 and 5). We confirmed the reading of the pressure of the South and North sensors of teh secondary pliers are good again (attachment 6). See pictures.